OpenStudy (anonymous):

solve the quadratic equation: 2x^2-7x+1=0

OpenStudy (anonymous):

and show work

OpenStudy (anonymous):

Help please

OpenStudy (anonymous):

According to the Quadratic Formula, x , the solution for Ax2+Bx+C = 0 , where A, B and C are numbers, often called coefficients, is given by : - B ± √ B2-4AC x = ———————— 2A In our case, A = 2 B = -7 C = 1 Accordingly, B2 - 4AC = 49 - 8 = 41 Applying the quadratic formula : 7 ± √ 41 x = ————— 4 √ 41 , rounded to 4 decimal digits, is 6.4031 So now we are looking at: x = ( 7 ± 6.403 ) / 4 Two real solutions: x =(7+√41)/4= 3.351 or: x =(7-√41)/4= 0.149

OpenStudy (anonymous):

thank you so so much!

OpenStudy (anonymous):

welcome